A sample of gas has a mass of 0.535 g . Its volume is 119 mL at a temperature of 80 ∘C and a pressure of 755 mmHg . Find the molar mass of the gas.
Answer #1

PV = nRT

where, P = pressure = 755 mmHg = 0.993 atm

V = volume = 0.119 L

n = number of moles

R = Gas constant

T = temperature = 80 + 273 = 353 K

0.993 * 0.119 = n * 0.0821 * 353

0.118 = n * 29.0

n = 0.118 / 29.0 = 0.00407 mole

number of moles = mass of gas / molar mass of gas

0.00407 mole = 0.535 g / molar mass of gas

molar mass of gas = 0.535 g / 0.00407 mole = 131 g/mol

Therefore, the molar mass of gas = 131 g/mol